OpenStudy (dimpalk104):

~~~Will medal and fan!!!~~~ What mass of water, in grams, would have to be decomposed to produce 13.9 L of molecular oxygen at STP? ???g

OpenStudy (photon336):

@dimpalk104 do you know the ideal gas equation pV =nRT?

OpenStudy (dimpalk104):

@Photon336 I know what ideal gas law is but there was no pressure or temperature given

OpenStudy (photon336):

@Dimpalk104 there is a trick here. STP means standard temperature and pressure. STP means 273K and 1 atm :)

OpenStudy (photon336):

we can easily find the number of moles of oxygen O2 from the ideal gas equation

OpenStudy (photon336):

\[\frac{ pV }{ RT } = n~number~moles \]

OpenStudy (photon336):

is this familiar @Dimpalk104 using this equation and the information given we can easily find the number of moles of oxygen O2

OpenStudy (dimpalk104):

@Photon336 okay so instead of using pV=nRT i used the 1mol/22.4L and found moles of oxygen and then using a balanced decomposition equation found the moles of H2O and then multiplied 18.02. i found the right answer. thank you for your help!

OpenStudy (photon336):

let's go back for a second

OpenStudy (photon336):

we need to use the ideal gas equation to find the number of moles of oxygen gas \[\frac{ 1~\cancel\atm*13.9~\cancel\L }{ 0.08~\cancel\L~mol^{-1}~\cancel\atm~\cancel\K^{-1}*273\cancel\K } = moles, O_{2}\]

OpenStudy (photon336):

\(\color{blue}{\text{Originally Posted by}}\) @Photon336 \[\frac{ pV }{ RT } = n~number~moles\] \(\color{blue}{\text{End of Quote}}\)

OpenStudy (photon336):

Then once we've found the number of moles of oxygen gas we need to go back to the balanced equation. here it is below.  once, we know this we can use the molar ratio to find the number of moles that we needed to react with this and then the number of grams.

OpenStudy (dimpalk104):

Yup thats what i did!! thanks again